Is bipolar junction transistor is a voltage controlled device
1
Expert's answer
2021-08-09T06:51:16-0400

Bipolar Junction Transistor (BJT) is a current controlled device as the collector's current is fully controlled by the base current.

The relationship between the current in the base, collector and emitter is given as

"I_e=I_b+I_c\\
I_c=\beta{I_b}\\
I_c=\alpha{I_e}"

Where "\alpha" and "\beta" are the gain of the Transistor in common emitter and common base mode respectively